Economists at the Wilson Company are interested in developing a production function for fertilizer plants. They have collected data on 15 different plants that produce fertilizer as shown in the table below.

Plant

Production

Capital

Labour

1

606

18891

700

2

566

19201

652

3

647

20655

823

4

524

15082

650

5

712

20300

859

6

488

16079

613

7

762

24194

851

8

443

11504

655

9

821

25970

901

10

398

10127

550

11

897

25622

842

12

359

12477

541

13

979

24002

949

14

332

8042

576

15

1065

23972

926

  1. Estimate the Cobb-Douglas production function Q = αLβ1Kβ2 where Q = output; L = labour input; K = capital input; and α, β1, and β2 are the parameters to be estimated. (Note: If the regression program on your computer does not have a logarithmic transformation, manually transform the preceding data into the logarithms before entering the data into the computer.)

  1. Test whether the coefficients of capital and labour are statistically significant.

  

  1. Determine the percentage of the variation in output that is "explained" by the regression equation.
  1. Determine the labour and capital production elasticities and give an economic interpretation of each value.

  1. Determine whether this production function exhibits increasing, decreasing, or constant returns to scale (ignore the issue of statistical significance).
Answer #1

a.

We know that the logarithmic model of Cobb-Douglas production function is given by:

lnQ = \alpha_0 + \beta_1lnL + \beta_2lnK + u

where α0 = ln(α)

So, calulating the log values of Q, L and K, we conduct the regression in Excel and obtain the results table as:

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.97353
R Square 0.947761
Adjusted R Square 0.939054
Standard Error 0.089983
Observations 15
ANOVA
df SS MS F Significance F
Regression 2 1.762813 0.881407 108.856 2.03E-08
Residual 12 0.097164 0.008097
Total 14 1.859977
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept -4.7488 0.809035 -5.86972 7.6E-05 -6.51154 -2.98607 -6.51154 -2.98607
lnK 0.414769 0.135089 3.070333 0.009711 0.120435 0.709103 0.120435 0.709103
lnL 1.077804 0.25046 4.3033 0.001026 0.532099 1.623509 0.532099 1.623509

So, the estimated double-log model is:

\widehat{lnQ} = -4.75 +1.08*lnL + 0.41*lnK

So, the estimated values for β1 and β2 are 1.08 and 0.41 respectively.

Since, α0 = ln(α), we can calculate the value of αas α = antilog(α0).

So, in this case,

\widehat{\alpha } = antilog(-4.75) = 0.0086

b.

By looking at the p-values in the regression table, we see that the p-value for the coefficient of lnK is 0.009. At 5% level of significance, since the p-value < 0.05, we reject the null hypothesis and infer that β2 is statistically significant.

The p-value for the coefficient of lnL is 0.001. At 5% level of significance, since the p-value < 0.05, we reject the null hypothesis and infer that β1 is statistically significant.

So, the coefficients of capital and labour are both statistically significant.

c.

The R-squared value is 0.947. This implies that 94.7% of the variation in output is explained by the regression equation.

d.

Since the above regression model is a double-log model, β1 and β2 represent elasticties of output with respect to labour and capital respectively.

\widehat{\beta_1 } = 1.08 represents the elasticity of output w.r.t labour.

This means that when the amount of labour increases by 1%, then on average, output increases by 1.08%, holding capital constant.

\widehat{\beta_2 } = 0.41 represents the elasticity of output w.r.t capital.

This means that when the amount of capital increases by 1%, then on average, output increases by 0.41%, holding labour constant.

e.

In order to find if the production function exhibits increasing, decreasing or constant returns to scale, we need to calculate the sum of estimated coefficients of labour and capital.

So, \widehat{\beta_1 } + \widehat{\beta_2} = 1.08 + 0.41 = 1.49

Since, this value is greater than 1, we can infer that the production function exhibits increasing returns to scale.